in a right-angled triangle one of the acute angles is 20 greater than then the other.fine the angles of the triangle
Found 2 solutions by solver91311, Cromlix:
Answer by solver91311(24713)



The sum of the measures of the interior angles of any triangle is 180 degrees. Since a right triangle has one angle that measures 90 degrees, the sum of the measures of the two acute angles must be 180 minus 90 equals 90 degrees.



Solve for

John

Egw to Beta kai to Sigma
My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism


Answer by Cromlix(4381)

Angles in triangle:
x, x + 20 and 90.
Therefore:
x + (x + 20) + 90 = 180
2x + 110 = 180
2x = 180 - 110
2x = 70
x = 35 degrees
Therefore angles are:
35 degs, 55 degs and 90 degs